# GRAph-Structured Pyramidal Whole Slide Image Representation

### v1

@@ 3.4 Inter-Magnification Convergence
$\lim_{m\to\infty}||h_{i}^{(3)}-h_{j}^{(3)}||_{2}=0;\lim_{m\to\infty}||h_{i}^{\prime(3)}-h_{j}^{\prime(3)}||_{2}=0;$ (5)
$and\lim_{m\to\infty}||h_{i}^{\prime\prime(3)}-h_{j}^{\prime\prime(3)}||_{2}=0$
Proof: Please see Inter-Magnification Convergence in Supplementary Materials.
Corollary 1
@@
Taking into account the fact that $h^{*}$, $h^{\prime*}$, and $h^{\prime\prime*}$ are not necessarily equal, our model is fusing node features in each magnification while it consults with other magnifications, and draw the conclusion via averaging nodes across three magnifications at the end of the convolutional layers by means of the readout module. This means that the final embedding of the graph is $\dfrac{h^{*}+h^{\prime*}+h^{\prime\prime*}}{3}$. We believe that this process helps the model reduce variance and uncertainty in making predictions as $m$ grows. To support this claim, we provide empirical evidence detailed in section Monte Carlo Test.
The structure of the graph has been designed in such a way that it does not get stuck in the bottleneck of over-smoothing, a common issue in deep GCNs [2]. Our intuition is that nodes in $\textbf{M}_{1}$, $\textbf{M}_{2}$, and $\textbf{M}_{3}$ are interacting via message passing, and the flow of intra-magnification information helps the model to keep its balance and continue the process of learning. Nevertheless, by increasing the number of GCN layers for the model to four or higher, the so-called over-smoothing problem will take place, which can possibly deteriorate the model’s performance. On the other hand, less than three layers of GCNs might not be able to fully capture the intra-magnification interactions. This leaves us with three layers of GCNs, which is equal to the graph’s diameter. In addition to this theoretical description, we empirically support our claim: please see Graph Depth in Supplementary Materials.
@@ 4.1 Data preparation
Table 1: Average performance on 3 folds and 10 random seeds
@@
Model params. inference Ovarian: five subtypes Bladder: two subtypes
Balanced Acc. F1 Score Balanced Acc. F1 Score
Trans-MIL $2.672M$ $0.053$ sec $0.647\pm 0.007$ $0.632\pm 0.005$ $0.868\pm 0.023$ $0.877\pm 0.013$
Ab-MIL $0.263M$ $0.005$ sec $0.692\pm 0.016$ $0.680\pm 0.014$ $0.919\pm 0.018$ $0.922\pm 0.016$
CLAM-SB $0.795M$ $0.009$ sec $0.627\pm 0.015$ $0.623\pm 0.010$ $0.908\pm 0.026$ $0.911\pm 0.023$
CLAM-MB $0.796M$ $0.012$ sec $0.620\pm 0.035$ $0.609\pm 0.030$ $0.901\pm 0.039$ $0.906\pm 0.037$
DGCN: latent $0.790M$ $0.018$ sec $0.654\pm 0.017$ $0.652\pm 0.024$ $0.835\pm 0.034$ $0.841\pm 0.035$
DGCN: spatial $0.790M$ $0.030$ sec $0.654\pm 0.009$ $0.652\pm 0.009$ $0.867\pm 0.015$ $0.875\pm 0.007$
PatchGCN: latent $1.385M$ $0.018$ sec $0.683\pm 0.003$ $0.675\pm 0.005$ $0.911\pm 0.031$ $0.919\pm 0.020$
PatchGCN: spatial $1.385M$ $0.032$ sec $0.672\pm 0.002$ $0.662\pm 0.005$ $0.896\pm 0.033$ $0.905\pm 0.021$
ZoomMIL $2.891M$ $0.036$ sec $0.701\pm 0.020$ $\mathbb{0.690}\pm 0.021$ $0.931\pm 0.008$ $0.933\pm 0.009$
GRASP (ours) $0.378M$ $0.005$ sec $\mathbb{0.727}\pm 0.036$ $0.689\pm 0.040$ $\mathbb{0.935}\pm 0.011$ $\mathbb{0.937}\pm 0.014$
@@
We use two datasets: Ovarian Carcinoma and Bladder Cancer, where the former consists of 948 WSIs with five histotypes including high-grade serous carcinoma (HGSC; 410), clear cell ovarian carcinoma (CCOC; 167), endometrioid carcinoma (ENOC; 237), low-grade serous carcinoma (LGSC; 69), and mucinous carcinoma (MUC; 65), and the latter consists of 262 WSIs with two histotypes including micropapillary (MicroP; 128) and conventional urothelial carcinomas (UCC; 134). A total of 1,133,388 patches of size $1000\times 1000$ pixels for the Ovarian dataset and 313,191 patches of size $1000\times 1000$ pixels for the Bladder dataset are extracted in multi-magnification mode (approximately 1.5 TB of Gigapixel WSIs). Patches being extracted such that they do not overlap at $\textbf{M}_{3}$ while overlapping at $\textbf{M}_{2}$ and $\textbf{M}_{1}$ is inevitable. From each magnification, $m\leq 400$ patches have been extracted per slide, as it’s been shown in [3, 27, 21] that a subset of patches is enough to represent WSIs. For each cancer dataset, we trained our proposed method in a 3-fold cross-validation and repeated the experiments ten times with different random seeds, where random seeds were randomly generated. In order to prevent data leakage in our cross-validation splits, we split the slides based on their patients, since some patients have more than one slide, meaning that slides were split in a way that all slides from the same patient remain in the same set. More information is available in Supplementary Materials.

### v2

@@ 3.4 Intra-Magnification Convergence
$\lim_{m\to\infty}||h_{i}^{(3)}-h_{j}^{(3)}||_{2}=0;\lim_{m\to\infty}||h_{i}^{\prime(3)}-h_{j}^{\prime(3)}||_{2}=0;\text{and}\lim_{m\to\infty}||h_{i}^{\prime\prime(3)}-h_{j}^{\prime\prime(3)}||_{2}=0.$ (5)
Proof: Please see Section 7.3 (Theoretical Analysis).
Note: It is worth mentioning that the assumptions made in Theorem 1 are minimal. In the case of $L_{2}$-bounded weights, it has been further explained in Wu et al. (2023); Cai & Wang (2020). The assumption that graph node features at $l=0$ are bounded is also minimal as we use a frozen features extractor, which ideally should not generate nondefinitive values for a finite feature vector.
With that, we show
$\left\|h_{i}^{(3)}-h_{j}^{(3)}\right\|_{2}\leq(\frac{1}{m+1})^{3}\left\|h_{i}^{\prime(0)}-h_{j}^{\prime(0)}\right\|_{2}\left\|W^{(2)}\right\|_{2}\left\|W^{(1)}\right\|_{2}\left\|W^{(0)}\right\|_{2},$
which essentially means that $\left\|h_{i}^{(3)}-h_{j}^{(3)}\right\|_{2}$ is upper bounded inversely with the reciprocal of $(m+1)^{3}$ Thus, as $m$ increases, the upper-bound gets tighter and eventually leads to $\lim_{m\to\infty}||h_{i}^{(3)}-h_{j}^{(3)}||_{2}=0$. As a result, we can conclude the following corollary.
Corollary 1.
@@
This is necessarily equivalent to pooling the nodes in magnification level, yet with a completely new approach than the traditional pooling layer, without further imposing computational load on the network for pooling. Taking into account the fact that $h^{*}$, $h^{\prime*}$, and $h^{\prime\prime*}$ are not necessarily equal, our model is fusing node features in each magnification while it consults with other magnifications, and draw the conclusion via averaging nodes across three magnifications at the end of the convolutional layers by means of the readout module. This means that the final embedding of the graph is $\dfrac{h^{*}+h^{\prime*}+h^{\prime\prime*}}{3}$. We believe that this process helps the model reduce variance and uncertainty in making predictions as $m$ grows. To support this claim, we provide empirical evidence detailed in Section 7.5.1 (Monte Carlo Test).
The structure of the graph has been designed in such a way that it does not get stuck in the bottleneck of over-smoothing, a common issue in deep GCNs Cai & Wang (2020). Our intuition is that nodes in $\textbf{M}_{1}$, $\textbf{M}_{2}$, and $\textbf{M}_{3}$ are interacting via message passing, and the flow of inter-magnification information helps the model to keep its balance and continue the process of learning. Nevertheless, by increasing the number of GCN layers for the model to four or higher, the so-called over-smoothing problem will take place, which can possibly deteriorate the model’s performance. On the other hand, less than three layers of GCNs might not be able to fully capture the inter-magnification interactions. This leaves us with three layers of GCNs, which is equal to the graph’s diameter. In addition to this theoretical description, we empirically support our claim in Section 7.5.5 (Graph Depth).
@@ 4.1 Data Preparation
We utilize three datasets: Esophageal Carcinoma (ESCA) from The Cancer Genome Atlas (TCGA), which includes 135 WSIs across two subtypes, and Ovarian Carcinoma and Bladder Cancer, where Ovarian Carcinoma consists of 948 WSIs with five histotypes, while Bladder Cancer contains 262 WSIs with two histotypes. These datasets were curated using HistoQC Janowczyk et al. (2019). A detailed breakdown of each dataset is available in Table 3.
4.2 Comparisons with State-of-the-Art
To compare with state-of-the-art approaches, we repeated these experiments with the same cross-validation folds and random seeds to have a fair comparison; the choice of ten random seeds is to capture statistical significance and reliability. For evaluating the models, we adopt Balanced Accuracy and F1 Score since these metrics show how reliable a model performs on imbalanced data, and more importantly on clinical applications. We compare our proposed model, GRASP, with models using different approaches to have a broad spectrum of evaluation. These models include Ab-MIL Ilse et al. (2018), Trans-MIL Shao et al. (2021), CLAM-SB Lu et al. (2021), and CLAM-MB Lu et al. (2021) from the attention/transformer-based family; ZoomMIL (2021) Thandiackal et al. (2022), H2MIL (2022) Hou et al. (2022), and HiGT (2023) Guo et al. (2023) from multi-magnification approaches since they have a hierarchical structure and are compatible with our patch extraction paradigm; and PatchGCN: latent & spatial Chen et al. (2021) and DGCN: latent & spatial Zheng et al. (2021) from graph-based learning approaches.
4.3 Subtype Prediction
Table 1 shows the comparison between our model and state-of-the-art methods based on Swin features, where GRASP outperforms all the competing methods on the Ovarian and Bladder datasets. Interestingly, Ab-MIL and CLAMs are the closest-performing methods to GRASP on these two datasets. On the ESCA dataset, however, ZoomMIL is the superior model with GRASP being the closest counterpart. Overall, GRASP is the superior model among all other models based on the average Balanced Accuracy on the three datasets.
Table 2 shows the comparison between our model and state-of-the-art methods based on KimiaNet features, where GRASP outperforms all the competing methods by a margin of $2.6\%-10.7\%$ Balanced Accuracy on the Ovarian dataset and $0.4\%-10.0\%$ on the Bladder dataset. It is worth mentioning that ZoomMIL is the closest-performing model to GRASP, although ZoomMIL has 7 times more parameters than GRASP. PatchGCN and DGCN are not performing comparably to GRASP, even though they are using spatial information that GRASP does not. This implies that a multi-magnification graph structure can potentially show more capability compared to other state-of-the-art approaches in terms of representing gigapixel WSIs. Moreover, single-magnification approaches are faster in terms of inference time than other approaches, especially CLAM-SB which has the lowest inference time. Inference times (per slide) have been calculated on the same machine for all models.
Table 1: The average performance on 3 folds and 10 random seeds based on Swin’s features. The best and second best average values are highlighted in bold and underlined, respectively.
@@ 4.3 Subtype Prediction
Model Params. Inference Ovarian: Five subtypes Bladder: Two subtypes ESCA: Two subtypes Average
Balanced Acc. F1 Score Balanced Acc. F1 Score Balanced Acc. F1 Score Balanced Acc.
Trans-MIL $2.672$M $0.019$ sec $0.297\pm 0.011$ $0.244\pm 0.011$ $0.830\pm 0.037$ $0.819\pm 0.030$ $0.626\pm 0.021$ $0.611\pm 0.021$ $0.584$
Ab-MIL $0.263$M $0.015$ sec ${0.643\pm 0.022}$ $0.647\pm 0.020$ $0.900\pm 0.013$ $0.884\pm 0.023$ $0.818\pm 0.010$ $0.812\pm 0.004$ $0.787$
CLAM-SB $0.795$M $0.014$ sec $0.546\pm 0.062$ $0.550\pm 0.065$ $\underline{0.903\pm 0.051}$ $\underline{0.902\pm 0.044}$ $\underline{0.877\pm 0.067}$ $0.861\pm 0.056$ $0.775$
CLAM-MB $0.796$M $0.015$ sec $0.558\pm 0.044$ $0.565\pm 0.042$ $\underline{0.903\pm 0.032}$ $0.901\pm 0.028$ $0.848\pm 0.055$ $0.833\pm 0.051$ $0.769$
DGCN: latent $0.790$M $0.098$ sec $0.224\pm 0.017$ $0.146\pm 0.017$ $0.725\pm 0.052$ $0.655\pm 0.108$ $0.763\pm 0.051$ $0.736\pm 0.059$ $0.570$
DGCN: spatial $0.790$M $0.086$ sec $0.210\pm 0.011$ $0.133\pm 0.012$ $0.700\pm 0.044$ $0.620\pm 0.074$ $0.660\pm 0.049$ $0.606\pm 0.053$ $0.523$
PatchGCN: latent $1.385$M $0.099$ sec $0.397\pm 0.039$ $0.362\pm 0.047$ $0.537\pm 0.011$ $0.351\pm 0.052$ $0.855\pm 0.076$ $0.847\pm 0.076$ $0.596$
PatchGCN: spatial $1.385$M $0.110$ sec $0.423\pm 0.042$ $0.390\pm 0.053$ $0.527\pm 0.020$ $0.336\pm 0.017$ $0.864\pm 0.080$ $0.859\pm 0.077$ $0.605$
ZoomMIL $2.891$M $0.024$ sec $0.640\pm 0.018$ ${0.648\pm 0.011}$ $0.899\pm 0.046$ $0.895\pm 0.037$ $\mathbb{0.889\pm 0.037}$ $\mathbb{0.895\pm 0.040}$ $\underline{0.809}$
HiGT $6.388$M $0.148$ sec $0.251\pm 0.037$ $0.184\pm 0.049$ $0.755\pm 0.041$ $0.717\pm 0.023$ $0.760\pm 0.050$ $0.744\pm 0.061$ $0.588$
H2MIL $0.829$M $0.092$ sec $\mathbb{0.671\pm 0.008}$ $\mathbb{0.667\pm 0.024}$ $0.900\pm 0.054$ $0.899\pm 0.044$ $0.854\pm 0.072$ $0.845\pm 0.084$ $0.808$
GRASP (ours) $0.378$M $0.024$ sec $\underline{0.669\pm 0.029}$ $\underline{0.654\pm 0.041}$ $\mathbb{0.905\pm 0.058}$ $\mathbb{0.906\pm 0.051}$ $\underline{0.877\pm 0.111}$ $\underline{0.872\pm 0.112}$ $\mathbb{0.817}$
Table 2: The average performance on 3 folds and 10 random seeds based on KimiaNet’s features. The best and second best average values are highlighted in bold and underlined, respectively.
Model Params. Inference Ovarian: Five Subtypes Bladder: Two Subtypes Model’s Average
Balanced Acc. F1 Score Balanced Acc. F1 Score Balanced Acc.
Trans-MIL $2.672M$ $0.019$ sec $0.647\pm 0.007$ $0.632\pm 0.005$ $0.868\pm 0.023$ $0.877\pm 0.013$ $0.758$
Ab-MIL $\mathbb{0.263M}$ $\underline{0.015}$ sec $0.692\pm 0.016$ $0.680\pm 0.014$ $0.919\pm 0.018$ $0.922\pm 0.016$ $0.806$
CLAM-SB $0.795M$ $\mathbb{0.014}$ sec $0.627\pm 0.015$ $0.623\pm 0.010$ $0.908\pm 0.026$ $0.911\pm 0.023$ $0.768$
CLAM-MB $0.796M$ $\underline{0.015}$ sec $0.620\pm 0.035$ $0.609\pm 0.030$ $0.901\pm 0.039$ $0.906\pm 0.037$ $0.761$
DGCN: latent $0.790M$ $0.098$ sec $0.654\pm 0.017$ $0.652\pm 0.024$ $0.835\pm 0.034$ $0.841\pm 0.035$ $0.745$
DGCN: spatial $0.790M$ $0.086$ sec $0.654\pm 0.009$ $0.652\pm 0.009$ $0.867\pm 0.015$ $0.875\pm 0.007$ $0.761$
PatchGCN: latent $1.385M$ $0.099$ sec $0.683\pm 0.003$ $0.675\pm 0.005$ $0.911\pm 0.031$ $0.919\pm 0.020$ $0.797$
PatchGCN: spatial $1.385M$ $0.110$ sec $0.672\pm 0.002$ $0.662\pm 0.005$ $0.896\pm 0.033$ $0.905\pm 0.021$ $0.784$
ZoomMIL $2.891M$ $0.024$ sec $\underline{0.701\pm 0.020}$ $\mathbb{0.690\pm 0.021}$ $\underline{0.931\pm 0.008}$ $\underline{0.933\pm 0.009}$ $\underline{0.816}$
HiGT $6.388$M $0.148$ sec $0.337\pm 0.044$ $0.288\pm 0.054$ $0.847\pm 0.067$ $0.842\pm 0.055$ $0.592$
H2MIL $0.829$M $0.092$ sec $0.653\pm 0.018$ $0.658\pm 0.032$ $0.876\pm 0.054$ $0.876\pm 0.048$ $0.764$
GRASP (ours) $\underline{0.378M}$ $0.024$ sec $\mathbb{0.727\pm 0.036}$ $\underline{0.689\pm 0.040}$ $\mathbb{0.935\pm 0.011}$ $\mathbf{0.937\pm 0.014}$ $\mathbf{0.831}$
Comparing Tables 1 and 2, all the models performed better with KimiaNet embeddings than with Swin embeddings, which is mostly because KimiaNet has domain knowledge and can provide more contextual features than Swin. Furthermore, GRASP, H2MIL, ZoomMIL, Ab-MIL, and CLAMs showcase robust generalization and effective performance even when utilizing features from different backbones, especially with GRASP being the most robust model.
Although Deng et al. (2023) has used attention score distribution to show their model is reliable across different magnifications, we want to step further and adopt a similar logic as first introduced in Selvaraju et al. (2017) to define the concept of energy of gradients for graph nodes (Please see Graph-Based Visualization 7.6 in Appendix). Therefore, for the first time in the field, we show that an AI model such as GRASP can learn the concept of magnification and behave according to the subtype and slide characteristics. To this end, we formulate an experiment to obtain a sense of each magnification’s influence on the model which leads to Figure 3. The main takeaway of this experiment is that depending on the subtype, the distribution of referenced magnifications by GRASP is different. From a pathological point of view, this finding fits our knowledge of the biological properties of each subtype. As an example, we conducted a case study on the bladder dataset, where the micropapillary subtype is known to be diagnosed generally in lower magnification owing to its morphological properties and the structure of micropapillary tumors, whereas UCC needs to be examined in higher magnifications due to its cell- and texture-dependent structure.
Furthermore, to examine whether GRASP understands the biological meaning of the data, i.e., differentiating between tumor vs non-tumor regions, we conduct a visualization experiment to plot the pixel-level heatmap of patches in multiple magnifications as depicted in Figure 4.
4.4 Ablation Study
Here, we design five experiments to evaluate our proposed model. Firstly, a Monte Carlo test on graph size, i.e., the number of nodes to investigate the impact of the number of nodes on the model’s performance (Section Monte Carlo Test 7.5.1). Secondly, analyzing model performance on individual or pairs of magnifications to study the effectiveness of multi-magnification representation (Section Magnification Test 7.5.2). Thirdly, we investigate the effect of different graph convolution types (Section Graph Convolutions 7.5.3). In the fourth experiment, we study how different models perform when all the patches from a WSI are used (Section Patch Number 7.5.4). Lastly, we empirically show that the graph depth of $d=3$ is the appropriate choice for our design (Section Graph Depth 7.5.5).
5 Conclusion
In this work, we developed GRASP, the first lightweight multi-magnification framework for processing gigapixel WSIs. GRASP is a fixed-structure model that can learn multi-magnification interactions in the data based on the idea of capturing both the inter- and intra-magnification information. This relies on the theoretical property of the model, where it benefits from intra-magnification convergence to pool the nodes rather than conventional pooling layers. GRASP, with its pre-defined fixed structure, has comparably fewer parameters than other state-of-the-art multi-magnification models in the field and outperforms the competing models in terms of average Balanced Accuracy over three complex cancer datasets using two different backbones. For the first time in the field, confirmed by two expert genitourinary pathologists, we showed that our model is dynamic in finding and consulting with different magnifications for subtyping two challenging cancers. We also evaluated the model’s decision-making to show that the model is learning semantics by highlighting tumorous regions in patches. Furthermore, we not only run extensive experiments to show the model’s reliability and stabilization in terms of its different hyperparameters, but we also provide the theoretical foundation of our work to shed light on the dynamics of GRASP in interacting with different nodes and magnifications in the graph. To conclude, we hope that the strong characteristics of GRASP and its straightforward structure, along with the theoretical basis, will encourage the modeling of lightweight structure-based design in the field of digital pathology for WSI representation.
6 MEANINGFULNESS STATEMENT
In digital pathology, a meaningful representation of life involves capturing the intricate, multi-scale structures of biological tissues, similar to how pathologists operate. GRASP (GRAph-Structured Pyramidal Whole Slide Image Representation) aids this process by modeling whole slide images as hierarchical graphs that integrate information across different microscopic magnification levels. This method, though lightweight, improves cancer subtyping accuracy and aligns computational analysis with human diagnostic processes, promoting deeper insights into tissue architecture and disease mechanisms.
7 Appendix
Figure 3: The histogram of consultations conducted by GRASP with different magnifications. First, this shows GRASP is actively dynamic in terms of capturing information from different magnifications benefiting from its multi-magnification structure. Second, information is distributed differently over magnifications depending on the subtype and slide, and there is no optimal magnification for a subtype. For example, in the Bladder dataset, ‘$(5x\&10x\&20x)$’ shows that the model needed to consult with all three magnifications for $19.3\%$ and $39.4\%$ of slides for MicroP and UCC, respectively; ‘$(5x)$’ shows that the model has mostly focused on only $5x$ magnification for $43.2\%$ and $5.6\%$ of slides for MicroP and UCC, respectively. This behavior is similar to pathologists, where they can diagnose massive MicroP tumors with lower magnifications, while they need to consult with higher magnifications to confirm a minuscule mass of MicroP tumors. On the other hand, UCC is hard to diagnose at lower magnifications and requires careful examination with different magnifications due to its morphological complexity, which fits the model behavior in proclivity to highlight more than one magnification for the majority of cases.
Figure 4: A case study on the Bladder dataset using KimiaNet features. a) Graph-based visualization: a random case from the subtype MicroP in the test data was selected to visualize its magnification heatmap where we show the absolute gradient in terms of each node. The $5x$ magnification contributes to $66.01\%$ of the whole energy model spent on this slide, meaning GRASP overall emphasizes more on $5x$ on this slide. Patch-based visualization: GRASP highlights patches of the three magnifications of a region of interest. In the second row, highlighted regions show the model has identified those areas as important while paying minimal attention to other regions. As confirmed by an expert pathologist, the model’s highlights on the three patches are tumors. The model can thus differentiate MicroP tumors from other tissue textures despite being trained for separating MicroP vs UCC. b) shows a similar case yet on the subtype UCC from a random slide in the test data. In this case, GRASP focuses on both $5x(47.45\%)$ and $10x(33.38\%)$ but is more interested in $5x$. As confirmed by the expert pathologist, the regions highlighted (yellowish areas in the second row) by the model are tumorous neighborhoods. Therefore, GRASP can differentiate UCC tumors from other textures and healthy cells across multiple magnifications.
7.1 Datasets
Table 3: Summary of the datasets used in this study.
Dataset Source No. of WSIs Histotypes/Subtypes
Ovarian Carcinoma Private Dataset 948 High-Grade Serous Carcinoma (HGSC): 410
Clear Cell Ovarian Carcinoma (CCOC): 167
Endometrioid Carcinoma (ENOC): 237
Low-Grade Serous Carcinoma (LGSC): 69
Mucinous Carcinoma (MUC): 65
Esophageal Carcinoma (ESCA) TCGA 135 Adenocarcinoma: 86
Squamous Cell Carcinoma: 49
Bladder Cancer Private Dataset 262 Micropapillary (MicroP): 128
Conventional Urothelial Carcinomas (UCC): 134
A total of 1,133,388 patches of size $1000\times 1000$ pixels for the Ovarian dataset, 602,874 patches of size $224\times 224$ from the ESCA dataset, and 313,191 patches of size $1000\times 1000$ pixels for the Bladder dataset are extracted in multi-magnification setting (approximately 2 TB of Gigapixel WSIs). Patches being extracted such that they do not overlap at $\textbf{M}_{3}$ while overlapping at $\textbf{M}_{2}$ and $\textbf{M}_{1}$ is inevitable. From each magnification, $m\leq 400$ patches (note that we use a large field of view, meaning this number eventually covers much of tissue regions) have been extracted per slide in both Ovarian and Bladder datasets, as it’s been shown in Chen et al. (2022a); Wang et al. (2023); Rasoolijaberi et al. (2022) that a subset of patches is enough to represent WSIs.
For patch-level feature extraction, we utilized two backbones: KimiaNet and Swin_base. Given that KimiaNet was trained on TCGA data in a supervised fashion, we intentionally refrained from extracting features from the ESCA dataset using this backbone to ensure an unbiased and leakage-free comparison. Conversely, we employed Swin, pre-trained on ImageNet, to extract features from all three datasets.
@@ 7.1 Datasets
For each cancer dataset, we trained our proposed method in a 3-fold cross-validation and repeated the experiments ten times with different random seeds, where random seeds were randomly generated, to ensure a rigorous comparison. In order to prevent data leakage in our cross-validation splits, we split the slides based on their patients, since some patients have more than one slide, meaning that slides were split in a way that all slides from the same patient remain in the same set.